OpenStudy (anonymous):

If the tan of angle x is 22 over 5, and the triangle was dilated to be two times as big as the original, what would be the value of the tan of x for the dilated triangle?

OpenStudy (anonymous):

The tan of the angle will be the same ...44/10 (Happens to be that the angle is about 77 degrees). tan (77 degrees approx) = 22/5 = 44/10= .....